Why does a perfect square result from adding two successive rows of even and odd integers? Example: 1+2 plus 1+2+3 = 9
1+2+3 plus 1+2+3+4 = 16

Answer by jim_thompson5910(35256)

Let n = last number in first row


If we generalize the problem, we get


1+2+3+...+n
1+2+3+...+n+(n+1)


Add the two rows to get
(1+2+3+...+n)+(1+2+3+...+n+(n+1))

Now simplify:

(1+1)+(2+2)+(3+3)+...+(n+n)+n+1

2(1)+2(2)+2(3)+...+2(n)+n+1

2(1+2+3+...+n)+n+1


Now use the formula 1+2+3+...+n = n(n+1)/2

2(n(n+1)/2)+n+1

n(n+1)+n+1

n^2+n+n+1

n^2+2n+1

(n+1)^2


So this shows that the sum of 1+2+3+...+n and 1+2+3+...+n+(n+1) is a perfect square and it shows us that the perfect square is (n+1)^2

In the case that n = 4, we get

(1+2+3+4)+(1+2+3+4+5) = 10+15 = 25

Take note that 25 = 5^2 = (4+1)^2 which confirms our answer.